# DISRetrieval: Harnessing Discourse Structure for Long Document Retrieval

### v1

@@ Abstract
Long document understanding has become increasingly crucial in natural language processing, with retrieval-based methods emerging as a promising solution to address the context length limitations of large language models (LLMs). However, existing approaches either treat documents as flat sequences or employ arbitrary chunking strategies, failing to capture the inherent discourse structure that guides human comprehension. We present DISRetrieval, a novel hierarchical retrieval framework that leverages linguistic discourse structure to enhance long document understanding. Our approach introduces three key innovations: (1) a discourse-aware document organization framework that utilizes rhetorical structure theory (RST) to create sentence-level hierarchical representations, preserving both semantic relationships and natural document flow; (2) an LLM-enhanced node representation technique that combines discourse structure with adaptive summarization to enrich tree nodes with contextual information; and (3) a hierarchical evidence retrieval mechanism that effectively selects relevant content while maintaining discourse coherence. Through comprehensive experiments on QASPER and QuALITY datasets, DISRetrieval demonstrates substantial improvements over existing methods in both token-level retrieval metrics and downstream question answering tasks. Our ablation studies confirm that incorporating discourse structure significantly enhances retrieval effectiveness across different document lengths and query types, validating the importance of linguistically-informed document representation in long-text understanding. Our code and datasets are publicly available at github/DreamH1gh/DISRetrieval to facilitate future research.
@@ 1 Introduction
Comprehending natural language documents is a fundamental task in natural language understanding [35]. With the increasing granularity and complexity of textual data, long document understanding has gained significant attention [3, 20]. Large language models (LLMs) have shown remarkable effectiveness for this task [10]. However, they still struggle with the challenges of memory constraints and restricted context lengths [4, 31]. Retrieval-based methods are a promising solution to these challenges [25, 17]. These methods first split long documents into small pieces and then use a retrieval model to select the most relevant evidence, thereby easing the burden of long-context reasoning for LLMs [42, 45]. The document splitting strategy is a critical factor that determines both the granularity of evidence selection and the effectiveness of the retrieval model.
The straightforward document splitting method is to segment a full document into a sequence of flattened pieces [24, 26]. However, it fails to capture the inherent structural relationships and semantic dependencies that exist within the long document, leading to suboptimal retrieval performance [29, 7]. To this end, several researchers have attempted various strategies to decompose long documents hierarchically, as illustrated in Figure 1. The most representative work is RAPTOR [34], which constructs a hierarchical tree for document organization through bottom-up semantic clustering. Additionally, an alternative bisection tree approach can be used, which only builds hierarchies between adjacent pieces to retain the original order of segmented pieces. Preliminary studies show that these tree-structure decomposition strategies are more effective for long document understanding.
@@
Figure 1: Comparison of document modeling approaches for long-text retrieval. Numbers (1-6) show sentence order in original document, with similar colors indicating semantic relationships. Four approaches are compared: (a) Flat sequential modeling, (b) Bottom-up semantic clustering of RAPTOR, (c) Bisection-based adjacent grouping, and (d) Our discourse-aware DISRetrieval that preserves both semantic and discourse structures.
Despite significant achievements in tree-based document organization, these methods generally ignore the real linguistically-informed discourse structures. Consider a document where the first sentence serves as a topic statement, followed by contrasting ideas (sentences 2-3), parallel supporting evidence (sentences 4-5), and a concluding statement (sentence 6). None of the previous retrieval-based methods have aligned with these types of knowledge structures. These discourse relationships naturally capture how humans organize and comprehend information in documents, making them particularly suitable for long document retrieval [15, 11, 36, 9]. This discourse-centric perspective not only aligns with human cognitive processes but also provides a more principled approach to document segmentation and representation compared to arbitrary chunking or purely semantic clustering methods [19, 14, 37].
In this work, we present the first approach that utilizes discourse information for retrieval-based long document understanding. Specifically, we propose DISRetrieval, which leverages rhetorical structure theory (RST) to organize long documents for retrieval. RST is a well-established discourse representation paradigm, using a hierarchical constituent tree to indicate the discourse structure of a document. The original RST exploits discourse-specific units as the bottom textual nodes. To better adapt our long document understanding, we ignore the inside-sentence discourse relationships, using only inter-sentence discourses. Thus, first we convert the original RST trees into sentence-level trees, and train an off-the-shelf RST model [41]. Second, we augment the non-leaf RST relation nodes with textual descriptions using an LLM agent. Finally, we construct an embedding tree based on the RST structure and employ it for query-based evidence retrieval to support LLM reasoning in long document understanding.
We conduct experiments on two benchmark datasets, QASPER [12] (a research paper QA dataset) and QuALITY [30] (a long-form reading comprehension dataset), to evaluate the effectiveness of our method. Results demonstrate that our approach significantly outperforms existing methods across various metrics. Specifically, our model achieves substantial improvements in both token-level retrieval metrics and downstream question answering tasks. Ablation studies validate the effectiveness of our discourse-aware tree construction and hierarchical retrieval mechanism, showing consistent gains over traditional chunking-based methods and flat retrieval approaches. These results demonstrate that our discourse-aware framework successfully captures both fine-grained semantic details and document-level patterns, while providing an effective solution for information propagation across different granularity levels in the hierarchy.
@@
  • We propose a novel discourse-aware hierarchical framework that leverages sentence-level discourse analysis to create semantically coherent document representations without relying on arbitrary chunking strategies.
  • We develop an innovative bottom-up node representation technique that effectively combines discourse structure information with LLM-based semantic enhancement.
  • Extensive experiments on two challenging datasets demonstrate significant improvements over existing methods, with up to 4% gains in retrieval metrics and consistent performance advantages across different model architectures.
@@ 2 Related Works
Document Comprehending.
Recent advances in natural language processing have revolutionized document processing and understanding [38, 43, 3]. While LLMs have demonstrated impressive capabilities, they face limitations due to context length constraints and memory bottlenecks [5, 10, 13, 1, 8]. Traditional approaches often struggle with long-form content, as they typically operate on short text segments without considering the broader document context [26, 16, 33]. Dense retrieval methods, particularly DPR [24], have shown promising results on passage-level tasks but encounter challenges when scaling to long documents where understanding the overall structure becomes crucial [25, 17]. Several innovative approaches have emerged to address these limitations - SEER focuses on extracting key information without conventional chunking [46], while CFIC explores chunking-free in-context retrieval [32]. These developments highlight the ongoing evolution in document comprehension strategies, particularly for handling extended textual content [39, 22, 42].
Retrieval-based Methods.
The retrieval-based paradigm has emerged as a promising solution for long document understanding, fundamentally operating by first segmenting documents into manageable pieces before retrieving the most relevant evidence for downstream tasks [25, 24, 17, 21]. While basic approaches rely on flat sequential splitting, this method fails to capture the inherent structural relationships and semantic dependencies within documents [18, 3, 43]. More sophisticated tree-based organization methods have been developed, with RAPTOR leading the way through its implementation of bottom-up semantic clustering [34]. Alternative approaches include bisection-based methods that maintain local coherence by grouping adjacent segments [44, 20]. Recent work has also explored hybrid hierarchical retrieval (HHR) and dense hierarchical retrieval (DHR), establishing connections between document-level and passage-level representations for more contextually informed retrieval [2, 27]. LongRefiner [23] converts retrieved documents into XML-format structure using a fine-tuned LLM to perform paragraph-level segmentation and top-down tree construction, but the length sensitivity inherent in this top-down approach creates challenges when structuring single, lengthy documents, despite its superior performance across multiple multi-document question-answering datasets. Meanwhile, these methods generally overlook the linguistic-informed discourse structures naturally present in documents [40, 28], leaving room for improvement in capturing document organization.
3 Methods
Retrieval-based methods have emerged as a promising solution for long document understanding, where the core idea is to first split documents into smaller pieces and then retrieve the most relevant evidence for downstream tasks. The most straightforward approach is to segment documents into flatten sequences. However, this approach cannot effectively capture the inherent structural relationships and semantic dependencies within long documents, resulting in suboptimal retrieval performance. Tree-based document organization methods provide enhanced document representation capabilities. Representative works like RAPTOR construct document hierarchies through bottom-up semantic clustering, while bisection-based approaches group adjacent segments to preserve local coherence. While these methods show improvements, they generally overlook the linguistic-informed discourse structures that naturally exist in documents.
We propose DISRetrieval, a discourse-aware retrieval framework that explores incorporating discourse structure into document retrieval. As a case study, we investigate the use of RST, a well-established discourse representation paradigm, to demonstrate how linguistic structures enhance document organization and retrieval effectiveness. By leveraging discourse relationships that reflect how humans organize and comprehend information, our framework maintains semantic coherence and structural integrity throughout the retrieval process, as illustrated in Figure 2. The algorithm of our approach is detailed in Algorithm 2 in Appendix C.
@@ 3.1 Discourse-Aware Tree Construction
3.1.1 RST Basics
RST provides a systematic framework for analyzing discourse structure in texts. An RST tree is a hierarchical representation where leaf nodes are text spans and internal nodes represent discourse relations between their children. These relations can be either mononuclear, where one span (the nucleus) is more central to the purpose of the writer than the other (the satellite), or multinuclear, where multiple spans are equally important. Common discourse relations include List, Contrast, Summary, and so on, which capture different ways that text segments relate to and support each other.
RST parsing is the task of automatically constructing RST trees from input text. Modern RST parsers typically employ neural networks to identify discourse boundaries and classify relations between text spans. These parsers first segment the text into elementary discourse units (EDUs), then incrementally build the tree structure by determining the relations between adjacent spans and combining them into larger discourse units. EDUs are the minimal building blocks in RST analysis, typically representing sub-sentence segments such as clauses. These fine-grained units serve as the leaf nodes of RST trees and form the basis for identifying discourse relationships in text.
3.1.2 Sentence-level Adaptation
While EDU-level discourse analysis provides fine-grained discourse structure, operating at the sentence level offers several advantages for long document understanding. First, sentences represent complete semantic units that maintain better coherence and interpretability compared to sub-sentence EDUs. Second, sentence-level segmentation significantly reduces computational complexity while preserving essential discourse relationships. Third, sentences align better with downstream retrieval tasks where the goal is to identify relevant text segments that can stand independently as evidence.
To enable sentence-level discourse analysis, we train a specialized discourse parser through the following process: (1) We merge EDUs belonging to the same sentence into single units from existing EDU-based discourse datasets. (2) For adjacent sentences, we determine their discourse relationship by identifying the lowest common parent node in the original EDU-based discourse tree. (3) We preserve these extracted relationships in a new tree structure with complete sentences as basic units. This conversion process allows us to train a parser $f_{\text{discourse}}$ that directly operates on sentence-level relationships while maintaining the essential discourse structure of texts.
Our core innovation lies in how we construct the document-level discourse tree through a two-phase iterative process that effectively preserves discourse relationships at multiple granularity levels:
@@ 3.1.2 Sentence-level Adaptation
where each tree captures how sentences connect and build upon each other within its respective paragraph, reflecting both semantic relationships and logical flow between adjacent sentences.
@@
The second phase obtains root text representation of paragraphs through bottom-up LLM enhancement. Since our discourse parser operates at sentence level, LLM enhancement helps simplify paragraph content into more manageable representations. For each internal node $v$ in the paragraph-level trees, we generate its text content according to:
$t_{v}=\begin{cases}f_{\text{LLM}}(t_{l},t_{r}),&\text{if }|t_{l}|+|t_{r}|\geqslant\tau\\ t_{l}\oplus t_{r},&\text{otherwise}\end{cases}$ (2)
where $t_{l}$ and $t_{r}$ are the text content of left and right children nodes. This adaptive summarization strategy reduces the complexity of multi-sentence content into more concise forms, making it more suitable for our sentence-level discourse parser. These root representations then serve as input for constructing the document-level discourse tree:
@@
As illustrated in Figure 3, this approach transforms rhetorical relations into concrete semantic summaries that effectively represent the content at different levels of the discourse hierarchy.
3.2 Node Representation
After obtaining both paragraph-level and document-level discourse trees from the previous construction process, we enhance the semantic representations of internal nodes through bottom-up LLM processing, integrate the enhanced trees into a unified hierarchical structure, and finally generate dense node embeddings to facilitate efficient semantic retrieval.
Input: Query $q$, discourse tree $T$, maximum evidence size $K$, leaf selection size $k$

### v2

# Beyond Chunking: Discourse-Aware Hierarchical Retrieval for Long Document Question Answering

@@ Abstract
Long document question answering systems typically process texts as flat sequences or use arbitrary segmentation, failing to capture discourse structures that guide human comprehension. We present a discourse-aware hierarchical framework that leverages rhetorical structure theory (RST) to enhance long document question answering. Our approach converts discourse trees into sentence-level representations and employs LLM-enhanced node representations to bridge structural and semantic information. The framework involves three key innovations: specialized discourse parsing for lengthy documents, LLM-based enhancement of discourse relation nodes, and structure-guided hierarchical retrieval. Comprehensive experiments on QASPER, QuALITY, and NarrativeQA demonstrate consistent improvements over existing approaches. Ablation studies confirm that incorporating discourse structure significantly enhances question answering across diverse document types.
Beyond Chunking: Discourse-Aware Hierarchical Retrieval for Long Document Question Answering
Huiyao Chen,   Yi Yang,   Yinghui Li,  Meishan Zhangthanks: Corresponding Author: Meishan Zhang,   Min Zhang Institute of Computing and Intelligence, Harbin Institute of Technology (Shenzhen), China {chenhy1018,yangi10010730,mason.zms}@gmail.com
@@ 1 Introduction
Document question answering represents a fundamental challenge in natural language processing, with applications spanning from academic research to enterprise knowledge management (Chen et al., 2017; Karpukhin et al., 2020). To date, large language models (LLMs) have achieved remarkable success on the task, particularly for short documents such as SQuAD (Rajpurkar et al., 2016), which has an average context length of 117 words (Reddy et al., 2019), reaching human-level performance with F1 scores exceeding 85% (Chowdhery et al., 2023; Malladi et al., 2023). However, as document length increases, their performance degrades significantly. For example, on challenging long document datasets such as QASPER, state-of-the-art LLM models only achieve performance with less than 50% F1 scores (Shaham et al., 2022; Bai et al., 2024). This performance gap highlights long document question answering as a critical and underexplored research frontier.
@@
Figure 1: Comparison of document modeling approaches for long document QA. Numbers (1-6) show sentence order in original document, with similar colors indicating semantic relationships. Four approaches are compared: (a) Flat sequential modeling, (b) Bottom-up semantic clustering, (c) Bisection-based adjacent grouping, and (d) Our discourse-aware approach that preserves both semantic and discourse structures.
The most straightforward approach to handling long documents involves flat sequential modeling, where texts are processed as linear sequences or divided into fixed-size chunks (Lewis et al., 2020; Guu et al., 2020a). This chunking-based paradigm has gained widespread adoption due to its simplicity and computational efficiency, enabling scalable processing of large document collections. Building upon this foundation, recent advances have introduced tree-based structures to better capture document organization and move beyond simple chunking. These innovations include semantic clustering methods like RAPTOR (Sarthi et al., 2024) that recursively group similar content, and bisection-based approaches that maintain local coherence through adjacent segment grouping (Liu and Lapata, 2019; Karpukhin et al., 2020). These hierarchical methods represent significant improvements by capturing semantic similarity and textual coherence, respectively, offering more sophisticated document modeling capabilities.
Discourse structure provides a principled linguistic foundation for document organization that goes beyond surface-level similarities. Consider a document where the first sentence serves as a topic statement, followed by contrasting ideas (sentences 2-3), parallel supporting evidence (sentences 4-5), and a concluding statement (sentence 6). These discourse relationships naturally capture how humans organize and comprehend information in documents, providing guidance for more semantically appropriate chunking and hierarchical organization in retrieval. As a mature theoretical framework, discourse analysis has demonstrated effectiveness across various NLP tasks (Feng and Hirst, 2014; Cohan et al., 2018; Søgaard et al., 2021), presenting an opportunity to develop more principled retrieval methods that align with human document comprehension patterns.
Therefore, in this work, we present DISRetrieval (DIScourse-aware hierarchical Retrieval), the first systematic approach that leverages rhetorical structure theory (RST) to enhance long document question answering through discourse-aware modeling. Our approach tackles three key technical challenges: First, we develop an effective adaptation strategy that enables traditional RST to work at appropriate granularity levels for computational efficiency while preserving semantic coherence. Second, we introduce LLM-based node enhancement that enriches intermediate nodes with both discourse structure and semantic content. Third, we design structure-guided evidence retrieval mechanisms that leverage discourse organization for more effective information extraction.
Comprehensive experiments across three challenging benchmarks: QASPER (Dasigi et al., 2021) for research paper understanding, QuALITY (Pang et al., 2022) for reading comprehension, and NarrativeQA Kočiský et al. (2018) for book-length narrative analysis demonstrate substantial improvements over existing methods. Detailed ablation studies validate the effectiveness of our discourse-aware modeling and structure-guided processing, showing consistent gains across diverse scenarios. Our framework successfully captures both fine-grained semantic details and document-level organizational patterns, providing a robust solution for discourse-informed question answering.
@@
  • A novel discourse-aware hierarchical framework that systematically incorporates RST for long document question answering.
  • An innovative LLM-enhanced hierarchical retrieval mechanism that enriches intermediate nodes with semantic content and enables multi-granularity retrieval.
  • Comprehensive empirical validation demonstrating consistent improvements across diverse datasets and model architectures.
Our code and model implementations will be publicly available to facilitate future research.
@@ 2 Related Works
Recent advances in natural language processing have achieved remarkable progress in document processing and understanding, with large language models demonstrating impressive capabilities across diverse text comprehension tasks (Brown et al., 2020; Chowdhery et al., 2023). However, these successes primarily focus on shorter texts, as extended documents present fundamental challenges due to context length constraints and computational complexity (Tay et al., 2021; Zaheer et al., 2020; Beltagy et al., 2020; Ding et al., 2023; Ainslie et al., 2023; Chen et al., 2024). Traditional approaches typically operate on short text segments without considering broader document context, leading to fragmented understanding (Liu and Lapata, 2019; Guo et al., 2022; Rae et al., 2020). Recent innovations such as SEER’s chunking-free key information extraction (Zhao et al., 2024) and CFIC’s in-context retrieval approach (Qian et al., 2024) highlight the ongoing evolution in document comprehension strategies for extended content (Wang et al., 2023; Izacard et al., 2023; Yu et al., 2023).
To tackle these challenges, retrieval-based methods have emerged as a promising paradigm, segmenting documents into manageable pieces before retrieving the most relevant evidence for downstream tasks (Lewis et al., 2020; Karpukhin et al., 2020; Guu et al., 2020a; Izacard and Grave, 2020). The evolution of retrieval-based approaches has progressed from simple flat segmentation strategies (Guu et al., 2020b) to sophisticated hierarchical organization methods. RAPTOR pioneered bottom-up semantic clustering for tree-based document representation (Sarthi et al., 2024), while bisection-based methods preserve local coherence through adjacent segment grouping (Zhang et al., 2020; Ivgi et al., 2023). Recent innovations include hybrid hierarchical retrieval systems that bridge document-level and passage-level representations (Liu et al., 2021; Arivazhagan et al., 2023), and LongRefiner’s XML-format structure conversion using fine-tuned LLMs for top-down tree construction, though its length sensitivity creates challenges for long documents (Jin et al., 2025). While these hierarchical approaches represent significant progress in document modeling, they primarily rely on surface-level semantic similarity or structural heuristics, overlooking the rich linguistically-informed discourse structures that naturally organize human-authored documents (Liu et al., 2019; Wolf et al., 2020).
3 Method
Discourse analysis provides systematic frameworks for understanding how textual segments relate beyond surface-level semantics, offering principled approaches to document organization Mann and Thompson (1988); Carlson et al. (2001). RST formalizes this understanding by representing documents as hierarchical trees where leaf nodes correspond to elementary discourse units (EDUs) and internal nodes encode rhetorical relations such as contrast, elaboration, and summary Marcu (2000). Recent advances in neural discourse parsing Yu et al. (2018); Kobayashi et al. (2020); Yu et al. (2022a) have enabled automatic RST construction at scale, facilitating its application across diverse NLP tasks Maekawa et al. (2024); Yuan et al. (2025).
To address the challenge of bridging between discourse theory and document retrieval, we propose DISRetrieval, a discourse-aware framework that systematically incorporates RST structure into document retrieval as shown in Figure 2. The complete algorithm is detailed in Algorithm 2 in Appendix A.
@@ 3.1 Discourse-Aware Tree Construction
3.1.1 Granularity-Adapted RST Processing
Traditional RST operates on fine-grained EDUs, which can create computational overhead and semantic fragmentation when processing long documents. We develop a granularity adaptation strategy that shifts RST processing to the sentence level, achieving better semantic coherence and efficiency while maintaining discourse interpretability. Our adaptation involves training a sentence-level discourse parser by systematically converting EDU-based datasets: we merge intra-sentence EDUs into coherent sentence units and determine inter-sentence relationships through lowest common ancestor analysis in the original EDU trees. This produces a specialized parser $f_{\text{discourse}}$ that operates directly on sentence-level discourse relationships, enabling efficient processing of long documents.
3.1.2 Complete Tree Construction Process
Our core innovation lies in developing a hierarchical discourse modeling framework that constructs complete document-level discourse trees through a systematic two-phase process, effectively bridging local sentence-level relationships with global document organization while preserving discourse coherence at multiple granularity levels.
@@ 3.1.2 Complete Tree Construction Process
where each tree $T_{i}$ captures how sentences connect and build upon each other within the paragraph, reflecting both semantic relationships and logical flow between adjacent sentences. This phase ensures that fine-grained intra-paragraph discourse structure is preserved.
@@
The second phase constructs the global document structure by first obtaining simplified representations of paragraphs through bottom-up LLM enhancement. For each internal node $v$ in paragraph-level trees, we apply adaptive processing:
$v^{*}=\begin{cases}f_{\text{LLM}}(v_{l},v_{r}),&\text{if }|v_{l}|+|v_{r}|\geqslant\tau\\ f_{\text{merge}}(v),&\text{otherwise}\end{cases}$ (2)
where $v_{l}$ and $v_{r}$ denote the left and right children of node $v$. The function $f_{\text{LLM}}$ generates concise summaries for larger text spans that align with our sentence-level discourse parser, while $f_{\text{merge}}(v)$ merges smaller, coherent subtrees into single leaf nodes to maintain semantic coherence.
The root representations obtained from each paragraph tree serve as input for constructing the document-level discourse tree:
@@
This approach transforms rhetorical relations into concrete semantic summaries that effectively represent the content at different levels of the discourse hierarchy (detailed illustration provided in Appendix A.3).
3.2 Discourse-Aware Node Representation
A critical challenge in discourse-based retrieval lies in bridging the semantic gap between abstract rhetorical relations and concrete textual content required for neural retrieval. We develop a semantic enhancement framework that transforms discourse structures into dense representations while preserving hierarchical properties through three key stages.
Bottom-up Semantic Enhancement.
The document-level discourse tree $T_{\text{doc}}$ initially contains only abstract structural relationships at internal nodes, lacking concrete semantic content necessary for retrieval. We apply bottom-up LLM enhancement following Equation 2 to generate meaningful textual representations for these structural nodes. This enhancement process systematically transforms each internal node from abstract structural placeholders into semantically rich representations that capture the combined content of their subtrees, ensuring all nodes possess concrete textual content suitable for downstream semantic matching operations.
Multi-Level Tree Integration.
We implement a substitution-based integration mechanism that combines the enhanced document-level tree $T_{\text{doc}}^{*}$ with paragraph-level trees $T_{i}$ into a unified discourse structure. Each leaf node in $T_{\text{doc}}^{*}$ is systematically replaced with its corresponding paragraph-level discourse tree:
$T_{\text{D}}=T_{\text{doc}}^{*}[l_{i}\leftarrow T_{i},\forall i\in\{1,2,...,n\}].$ (4)
This integration enables multi-granularity discourse modeling, where fine-grained intra-paragraph sentence relationships coexist with coarse-grained inter-paragraph discourse connections within a single hierarchical structure.
Node Encoding.
To enable efficient retrieval, we transform the integrated discourse tree $T_{\text{D}}$ into dense embeddings using a pre-trained encoder:
$\mathbf{e}_{v}=f_{\text{enc}}(v),\quad\forall v\in\mathcal{N}(T_{\text{D}}).$ (5)
The resulting embedding tree $\mathcal{T}_{\text{emb}}=(T_{\text{D}},\{\mathbf{e}_{v}\})$ preserves both the hierarchical discourse structure and semantic information, enabling structure-aware retrieval through dense vector operations.
3.3 Structure-Guided Evidence Retrieval
We introduce a structure-aware retrieval mechanism that leverages the hierarchical nature of discourse trees to select evidence at multiple granularity levels. Unlike flat retrieval methods that operate on uniform text segments, our approach exploits discourse structure to implement a dual-selection strategy that balances local semantic relevance with global discourse coherence.
Retrieval Process.
Given a query $q$, we first compute semantic similarities for all nodes in the embedding tree:
$\text{score}(v)=\text{cos}(f_{\text{enc}}(q),\mathbf{e}_{v}),\forall v\in\mathcal{N}(T_{\text{D}}).$ (6)
Dual-Selection Strategy.
Nodes are ranked by relevance and processed through two complementary selection mechanisms:
Direct leaf selection: High-relevance sentences selected immediately Hierarchical expansion: Internal nodes trigger top-$k$ subtree selection Redundancy elimination: Automatic tracking prevents duplicate selection
For leaf nodes with high relevance, we select them directly. For internal nodes, we perform controlled subtree expansion by selecting the top-$k$ most relevant unused leaves from their subtrees. This ensures both specific sentence-level evidence and coherent discourse segments are captured while eliminating redundancy.
The complete algorithm is detailed in Algorithm 1, which naturally balances local semantic relevance with global discourse coherence, enabling retrieval at appropriate granularity levels determined by the document’s discourse structure.
Input: Query $q$, discourse tree $T$, evidence limit $K$, subtree selection size $k$